# Generalization v.s. Memorization: Tracing Language Models’ Capabilities Back to Pretraining Data

### v2

@@ 8 Conclusion
In this paper, we introduce the task-gram language model, a scalable approach for modeling the task-relevant distribution of pretraining data. We trace the capabilities of LLMs back to this data by defining distributional memorization, measured through the Spearman correlation between task-gram LM probabilities and LLM probabilities. Through extensive experiments using Pythia models across three distinct tasks, we find that LLMs tend to memorize more when engaged in simpler, knowledge-intensive tasks, while they generalize more in harder, reasoning-intensive tasks. Our analysis provides a comprehensive examination of the origins of LLM capabilities and offers a scalable framework for investigating the fine-grained task-relevant characteristics of pretraining corpora.
@@ Appendix A Limitations
While this paper provides valuable insights into how large-scale pretraining corpus contributes to the emergent abilities of LLMs through $n$-gram search, there are a few limitations that we want to list out. First, the model we use, Pythia [Biderman et al., 2023], and the pretraining corpus, Pile [Gao et al., 2020], are slightly outdated and have been outperformed by many new open-source LLMs. However, most open-source LLMs lack corresponding pretraining data and have limited model sizes and pre-training checkpoints, hindering scaling effect studies. The current WIMBD system also has limitations in searching larger corpora like Dolma [Soldaini et al., 2024] (3T tokens), calling for improved searching and retrieval methods. The quality of task-relevant $n$-gram pairs is highly sensitive to the filtering method, and while the current embedding similarity-based approach is effective, better filtering methods could significantly enhance the analysis, which is left for future research.
@@ Appendix B Ethics Statement
Appendix C Experiment Details
@@
Appendix D $n$-gram Pair Examples
@@
Appendix E Alternative methods for Identifying $n$-gram Pairs
In order to isolate the best way to identify $n$-gram pairs, we tried a few alternative methods that did not yield high-quality alignment. More specifically, we used both LLMs and named-entity-recognition (NER) models to classify pretraining data chunks as ”containing $n$-gram pairs.” These methods ultimately failed, as we were not able to identify $n$-gram pairs from any other text with consistency, with the results having too many false positives and false negatives. In contrast, the method we employed in the paper, of first searching for $n$-grams and then filtering for semantic similarity, yielded much more consistent results.
Figure 9: Comparison of model distribution probabilities assigned to text chunks. Orange (unsupervised): No pairs present. Blue (supervised) Pairs actually present.
Figure 10: Count of examples assigned as Yes or No across different probability ranges for P(yes) in pretraining data. Blue (Yes): Pair is contained. Red (No): No pair. The labels were assigned by human evaluators
Figure 11: Assigned Probability of Model to different chunks containing $n$-gram Pairs for various tasks in the FLAN family of tasks. ”No Pairs” class is used as a control.
Appendix F Model Generations vs. $n$-gram Frequency in Pretraining

### v3

@@ 9 Conclusion
In this paper, we introduce the task-gram language model, a scalable approach for modeling the task-relevant distribution of pretraining data. We trace the capabilities of LLMs back to this data by defining distributional memorization, measured through the Spearman correlation between task-gram LM probabilities and LLM probabilities. Through extensive experiments using Pythia models across four distinct tasks, we find that LLMs tend to memorize more when engaged in simpler, knowledge-intensive tasks, while they generalize more in harder, reasoning-intensive tasks. Our analysis provides a comprehensive examination of the origins of LLM capabilities and offers a scalable framework for investigating the fine-grained task-relevant characteristics of pretraining corpora.
@@ Appendix A Limitations
While this paper provides valuable insights into how large-scale pretraining corpus contributes to the emergent abilities of LLMs through $n$-gram search, there are a few limitations that we want to list out. First, the model we use, Pythia [Biderman et al., 2023], and the pretraining corpus, Pile [Gao et al., 2020], are slightly outdated and have been outperformed by many new open-source LLMs. However, most open-source LLMs lack corresponding pretraining data and have limited model sizes and pre-training checkpoints, hindering scaling effect studies. The current WIMBD system also has limitations in searching larger corpora like Dolma [Soldaini et al., 2024a] (3T tokens), calling for improved searching and retrieval methods. The quality of task-relevant $n$-gram pairs is highly sensitive to the filtering method, and while the current embedding similarity-based approach is effective, better filtering methods could significantly enhance the analysis, which is left for future research.
@@ Appendix B Ethics Statement
Appendix C Related Work Details
Understanding LLMs’ capabilities from training data Prystawski et al. [2023] and Wang et al. [2024] discuss how the reasoning ability of language models is a consequence of their pretraining data. Prystawski et al. [2023] discusses how chain-of-thought reasoning is effective in autoregressive language models because of local structure within pretraining data, and Wang et al. [2024] derives novel conclusions from known facts by aggregating reasoning paths seen in pretraining data. On the other hand, Xie et al. [2022] and Wang et al. [2023] discuss how in-context learning is a by-product of learning the pretraining data distribution. They both suggest that language models learn to implicitly infer a latent variable from the given prompt, as the pretraining data is generated from some unknown latent variable. Additionally, Chan et al. [2022] propose that the distributional properties of training data drive emergent in-context learning behaviors in large language models, whereas Razeghi et al. [2023] shows the influence the pretraining data on the mathematical abilities of LLMs. Chen et al. [2024] also highlights the significance of parallel structures in pretraining data for the emergence of in-context learning.
However, the small-scale nature of such analysis is antithetical to the commonly believed main driving factor behind the performance of LLMs: scaling. Recently, Kirchenbauer et al. [2024] proposes to provide statistical evidence of the dependence of a target model capabilities on subsets of its training data, by estimating the data distribution with an embedding-induced kernel. However, their estimation is based on a very small portion of the pretraining data (around 0.3%) as computing the embeddings of a huge dataset is very non-trivial. To get a better estimation of the whole distribution of the pretraining data, Elazar et al. [2024] construct a retrieval system, WIMBD, that can efficiently search $n$-gram phrases over hundreds and thousands of GBs of pretraining data. Merrill et al. [2024] propose an efficient data structure that enables unbounded-length $n$-gram searches in massive pretraining datasets, and find that larger LLMs generate less novel large $n$-grams compared to human written texts, and [Shaib et al., 2024] shows how syntactic templates from the training data causes models to re-use such templates after training. Liu et al. [2024] proposes to build a prefix-based efficient retrieval system for pretraining corpora, and then construct large-scale $\infty$-gram language models to estimate the distribution of text corpora. However, such distribution only models local contextual dependency, which might not be useful for understanding complex LLM capabilities.
Memorization v.s. generalization The phenomenon of machine learning models being able to perfectly memorize the training data has been studied in many previous works. Most of them define LLM memorization as exactly recalling the training examples by designed prompting, including the memorization of rare long-tail data, like private information [Zhang et al., 2023], and the contamination of testing sets [Jiang et al., 2024]. Carlini et al. [2022] found that the exact copy and pasting behaviors are more prevalent in larger LMs. Hartmann et al. [2023] provides a comprehensive summarization of LLM memorization.
Several papers have studied the interplay between memorization and generalization of training data. Feldman [2020] prove that memorizing the training data is in fact required for optimal generalization on testing data. Works along this line [Feldman and Zhang, 2020, Zhang et al., 2023] extend the original definition of memorization by quantifying the extent of memorizing a training example with the performance difference when including and excluding this specific example in training data. However, this definition is impractical for large-scale analysis of LLMs as it requires retraining an LLM from scratch to analyze one data point. In this paper, we propose a new definition of distributional memorization by using $n$-gram counts, which is more suitable for large-scale analysis with LLMs.
Appendix D Experiment Details
Dataset choice
The choice of dataset comes from the consideration of balancing different types of tasks, and we believe the combination of translation + factual QA + reasoning can well represent the spectrum of possible tasks. To enhance the selection of reasoning tasks, we also add GSM8K as a new task as described in point 1 of the general response. We found both the performance v.s. n-gram count results and the distributional generalization results on GSM8K align with our previous findings.
Embedding model choice
The choice of the embedding model comes from accommodating the need for different tasks. We use the LASER embedding model for the Translation task because it is specialized for translation. We use E5 for other tasks because it is a general-purpose sentence embedding model that performs cosine-similarity-based retrieval.
Task-gram table construction
For WMT, we use the Europarl [Koehn, 2005] parallel corpus to construct a large task-gram table. For TriviaQA, we use its training set, while for MMLU, we use the testing set as there is no training set available. We employ two different models to filter the $n$-grams pairs. For translation tasks, we use the LASER embeddings [Schwenk and Douze, 2017], which provide language-agnostic sentence representations, in order to assess cross-lingual similarity. We mine $n$-gram pairs from the Europarl corpus, which consists of around 20 million parallel sentences extracted from the proceedings of the European Parliament [Koehn, 2005]. For all other tasks, we use E5 [Wang et al., 2022], which are multilingual sentence representations trained using contrastive learning on a diverse range of tasks. For TriviaQA, we mine the $n$-gram pairs from TriviaQA’s training set. For MMLU, since the training set is very small ($100-500$ examples for each task), we mine the $n$-gram pairs from the test sets directly.
In WMT, the cosine similarity thresholds we use are: 0.85, 0.8, 0.75, and 0.7 for 2 to 5-gram pairs, respectively; For TriviaQA and MMLU, the values are 0.75 and 0.65 for 3 and 5-gram pairs, respectively. We use lower thresholds for larger $n$-grams because larger $n$-grams inherently impose stricter alignment, and are therefore less likely. We show a cosine similarity sensitivity ablation in Figure 6.
Figure 6: Visualization of distributional memorization with different-sized Pythia models on TriviaQA with different cosine similarity threshold $\gamma\in\{0.7,0.75,0.8\}$. The trend of distributional memorization does not change with different thresholds.
@@ Appendix D Experiment Details
Figure 7: GSM8K accuracy v.s. $n$-gram pair count in Dolma with different OLMo model sizes.
Appendix E Prompt Optimization Details
Dataset Memorization Generalization
TriviaQA Deliver an exact single word or concise phrase in response to the factual question. (avg. 3-gram count: 557948.5) Formulate a distinctive and concise term or phrase to clearly answer the factual question. (avg. 3-gram count: 1714.9)
GSM8K Carefully analyze each math word problem presented, break it down step-by-step, and clearly state the final answer. (avg. 3-gram count: 45028.9) Dissect each math word problem into straightforward, logical steps; solve each part systematically for precise solutions. (avg. 3-gram count: 43.1)
Table 2: Prompts optimized for memorization and generalization for TriviaQA and GSM8K.
=Text in the square brackets are comments that are not involved in the actual prompt.
Meta prompt for prompt optimization **Task Description**: You are tasked with optimizing a given prompt to guide an open-source language model (LM) in completing a specific task effectively. You will receive: - The current prompt for the task. - Its corresponding memorization score (Average frequency of task-related n-grams found in the LMs pretraining corpus). - A few example input-output pairs illustrating the intended task. - A history of previous prompt optimization iterations. **Optimization Goals**: Clearly describe the intended task with a general instruction that effectively guides the LM to perform the task. [If trying to encourage reasoning …] Minimizing the memorization score of the updated prompt. The memorization score reflects the distributional correlation between the prompt and the LMs pretraining corpus. A lower score encourages the LM to generate more novel outputs. [If trying to encourage memorization …] Maximize the memorization score of the updated prompt. The memorization score reflects the distributional correlation between the prompt and the LMs pretraining corpus. A higher score encourages better alignment with the LMs learned knowledge. **Example task input-output pairs**: [Example input-output pairs for current task] **Output**: Produce an updated prompt that balances clarity of task instruction with an lower memorization score.
Appendix F $n$-gram Pair Examples
@@
Appendix G Model Generations vs. $n$-gram Frequency in Pretraining